Assume that you invest $100 at time zero and that your investment is worth $110 one period later.  A.  Did you make any money? How much in dollars?

How much in percent return on investment?  B. Use the rate of return on investment from your previous answer as the discount rate and compute what the NPV of your investment was.  What is the NPV?  Did you make any money?  C.  What does this say about a zero-NPV investment?
Answer #1

a: Amount of money made = $110-100=$10

b: Rate of return on investment = 10/100 = 10%

NPV =C0+ CF1/(1+r)^1

= 110/1.1^1 -100

= 0

Since NPV is zero, no money was made.

c: Zero NPV investment implies that the rate of return is the IRR of the investment which is 10%. Also, the rate of return equals the opportunity cost of funds and hence the NPV is zero.